OpenStudy (anonymous):

The difference of a number and 9 is equal to the quotient of 6 and 3. Find the number.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

n-9=2 n=11

OpenStudy (angela210793):

x-9=6/3=2 x=2+9 x=11

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

BECAUSE IM RESOLVIN THEM BUT IM JUST PUTTIN THEM HERE TO MAKE SURE IM DOING IT RIGHT